From initial conditions to galaxies

Corentin Cadiou March 29, 2018

1/100 2/100

Introduction

3/100

Classifying galaxies

4/100

A few parameters:

1. M_{200}

2. color / SFR

3. elliptical / spiral ?

5/100

Dependance of the parameters

• On space – image of H-AGN

• On z – show redshift evolution / ref

6/100

What to do?

Fact A  Fact B

7/100